# Characterisation of the big Ramsey degrees of the generic partial order

### v4

@@ Acknowledgement
We would like to thank the two anonymous referees for remarks which significantly improved clarity of this paper, and to Štěpán Vodseďálek for pointing out an important typo in Definition 1.5.
D. Ch., J. H. and M. K. were supported by the project 21-10775S of the Czech Science Foundation (GAČR). This article is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 810115). M. B. and J. H. were supported by the Center for Foundations of Modern Computer Science (Charles University project UNCE/SCI/004). N. D. was supported by National Science Foundation grant DMS-1901753. L. V. was supported by Beatriu de Pinós BP2018, funded by the AGAUR (Government of Catalonia) and by the Horizon 2020 programme No 801370. A. Z. was supported by NSF grant DMS-2054302 and NSERC grants RGPIN-2023-03269 and DGECR-2023-00412.

### v5

@@ Acknowledgement
We would like to thank the two anonymous referees for remarks which significantly improved clarity of this paper, and to Štěpán Vodseďálek for pointing out an important typo in Definitions 1.5 and 9.6.
D. Ch., J. H. and M. K. were supported by the project 21-10775S of the Czech Science Foundation (GAČR). During the revisions D. Ch., and J. H. were supported by project 25-15571S of the Czech Science Foundation (GAČR). This article is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 810115). M. B. was supported by the Center for Foundations of Modern Computer Science (Charles University project UNCE/SCI/004). N. D. was supported by National Science Foundation grants DMS-1901753 and DMS-2300896. L. V. was supported by Beatriu de Pinós BP2018, funded by the AGAUR (Government of Catalonia) and by the Horizon 2020 programme No 801370. A. Z. was supported by NSF grant DMS-2054302 and NSERC grants RGPIN-2023-03269 and DGECR-2023-00412.